What is the binomial expansion of a^n -b^n. . ...... . .

Grade:12

1 Answers

Samyak Jain
5 years ago
a- bn  =  a[1 – (b/a)]
(b/a) is the common ratio r of the finite geometric series :
1 + r + r+ r3 + ….............. + rn-1 , where r = (b/a)
1 + r + r+ r3 + ….............. + rn-1 = (1 – rn)/ (1 – r)
(1 – rn) = (1 + r + r+ r3 + ….............. + rn-1) (1 – r)
i.e. 1 – (b/a) = {1 + (b/a) +(b/a)+ (b/a)3 + ….............. + (b/a)n-1} {1 – (b/a)}
      1 – (b/a) = {1 + (b/a) +(b/a)+ (b/a)3 + ….............. + (b/a)n-1} {a – b} / a
Therefore, a- bn  =  an {1 + (b/a) +(b/a)+ (b/a)3 + ….............. + (b/a)n-1} {a – b} / a
  a- bn  =  an-1 {1 + (b/a) +(b/a)+ (b/a)3 + ….............. + (b/a)n-1} {a – b}
              = (a – b) (an-1 + an-2 b + an-3 b+ a bn-2 + bn-1)